### v1

# Chameleon Channels: Measuring YouTube Accounts Repurposed for Deception and Profit

Abstract
Online content creators spend significant time and effort building their user base through a long, often arduous process, which requires finding the right “niche” to cater to. So, what incentive is there for an established content creator known for cat memes to completely reinvent their page channel and start promoting cryptocurrency services or cover electoral news events? And, if they do, do their existing subscribers not notice?
We explore this problem of repurposed channels, whereby a channel changes its identity and contents. We first characterize a market for “second-hand” social media accounts, which recorded sales exceeding USD 1M during our 6-month observation period. By observing YouTube channels (re)sold over these 6 months, we find that a substantial number (37%) are used to disseminate potentially harmful content, often without facing any penalty. Even more surprisingly, these channels seem to gain rather than lose subscribers.
To estimate the prevalence of channel repurposing “in the wild,” we also collect two snapshots of 1.4M quasi-randomly sampled YouTube accounts. In a 3-month period, we estimate that $\sim$0.25channels—collectively holding $\sim$44M subscribers—were repurposed. Through a set of experiments, we confirm that these repurposed channels share several characteristics with sold channels—mainly, the fact that they had a significantly high presence of potentially problematic content. Across repurposed channels, we find channels that became disinformation channels, as well as channels that link to web pages with financial scams. We reason that abusing the residual user trust placed on these channels is advantageous to financially- and ideologically-motivated adversaries. This phenomenon is not exclusive to YouTube and we posit that the market for cultivating organic audiences is set to grow, particularly if it remains unchallenged by mitigations, technical or otherwise.
1 Introduction
Online content creators and influencers are central to digital spaces. Teens and adults in the United States alike consume hours of content every day from their favorite creators [1, 2]: from news [3], and recommendations for products and services [4] to advice on financial, professional, and medical topics [5]. Being a content creator is a highly sought after profession, with more than 57% GenZers saying they would like to become an influencer if given the opportunity [6]. In that context, it is unsurprising that tools and services dedicated to growing one’s online influence abound. Offerings range from cheap bulk bot-based engagement to increasingly more sophisticated products involving engagement from real humans.
A new product type has gained popularity in recent years: the sale of credentials to an account with a pre-established audience and reputation. A recent study found over 38,000 accounts and an advertised value of over USD 64 million [7]. According to Beluri et al. [7], Fameswap boasts the highest number of sellers and has experienced an almost 9x growth since being first described by Chu et al. in 2022 [8].
A key question is whether platforms like Fameswap constitute a new paradigm for online engagement—as opposed to just a new packaging for already known artificial engagement strategies. If these ready-made accounts represent a new type of product, what are the implications for the platforms whose accounts are commercialized? We aim to answer these questions by observing how these accounts are used after being sold. We conduct this observation on YouTube by leveraging the persistent nature of channel IDs—invariant identifiers that map to a channel, and do not change even if the channel entirely alters its identity by changing handles, title, description, and/or content–––what we call repurposing a channel, shown in Figure 1.
We hypothesize that repurposing channels alleviates a critical constraint for financially and ideologically motivated adversaries: rapid access to distribution without the resource expenditure required to cultivate an audience from scratch. To evaluate this hypothesis, we explore the following research questions:
  • RQ1: How are channels sold in online marketplaces repurposed?
  • RQ2: Do repurposed channels disseminate potentially problematic content?
  • RQ3: To which extent are channels repurposed on YouTube?
  • RQ4: What are indicators of channel reuse?
To investigate these questions, we conducted a mixed-methods study in which we characterize the Fameswap marketplace and collect YouTube channels advertised for sale, exploring the time it takes until a channel is repurposed, the change in subscribers after the change, and the rate of repurposing. We then collect a large quasi-random sample of YouTube channels from an analytics platform named Social Blade, and identify repurposed channels in the wild. Using our Fameswap and Social Blade samples of repurposed channels, we annotated each channel based on potentially problematic content categories.
Our work provides an in-depth view of an emerging vector that currently faces few mitigations. Its implications are the same for all other platforms for which there are accounts for sale.
2 Related Work
There is considerable related work in areas adjacent to this study, which we will discuss next. In particular, we explain how our study differs from these previous efforts.
Abusing Reputation Systems. Online reputation refers to metrics that signal user trustworthiness. First developed for online marketplaces, reputation systems—reviews, ratings, etc.—were key to their success [9]. These systems were later adopted by social media to incentivize contributions [10], encourage participation [11, 12], and signal authenticity [13]. Today, they take the form of likes, followers, badges, and similar indicators.
These indicators are valuable targets for manipulation. Over the past two decades, researchers have documented widespread abuse across platforms [14, 15, 16, 17, 18]. Misuse has evolved from bots and compromised accounts [19, 20, 21, 22] to more human-like tactics: crowdsourced engagement [23, 24, 25], automation of real user accounts [26], collusion rings [27, 28], and click farms [29, 30].
Account hijacking has similarly shifted toward high-value targets. Rather than compromising many small accounts, actors now focus on fewer but influential ones [31]. This shift is tied to the ease of monetizing influence, particularly in finance and crypto spaces. Notable examples include the SEC’s hijacked account used for market-moving tweets [32] and Vitalik Buterin’s account used to spread phishing links [33].
Reputation abuse also occurs without hijacking. Liu et al. [34] show that scammers use legitimate-looking YouTube channels (e.g., “Ripple XRP” with 114,000 subscribers) to host ephemeral livestreams and direct viewers to off-site scams.
Licit Yet Harmful Online Influencers. Influential accounts that promote potentially harmful content but within gray areas have become an increasingly prevalent issue. For example, many reported cases of influencers promoting “meme coins” have left viewers with little recourse after the coin collapses [35, 36]. It is unclear whether YouTube’s terms-of-service restrict this behavior, whose (il)legality is not even completely settled. Similarly, influential accounts have also been used to disseminate problematic narratives, such as smear campaigns, before elections. The latest case occurred just last year, when a criminal investigation linked payments from a Russian disinformation campaign to American influencers [37]. This example also illustrates the increasing interplay between financially motivated actors (in this case, influencers) and ideologically motivated actors (state sponsors).
Disinformation and Underground Markets. Services to assist or implement disinformation campaigns have often overlapped with underground markets and communities [38, 39, 40, 41]. On the one hand, these marketplaces specialize in the dissemination of content that is frequently at odds with platform guidelines, for example, advertising restricted products (e.g., pharmaceuticals, tobacco products, unregulated gambling websites, etc.), including disinformation. On the other hand, these marketplaces also provide tools and infrastructure that can aid disinformation operators (e.g., coordinated engagement, bulletproof hosting, DDoS protection, advertising networks, etc.) [42]. Most importantly, off-platform sources of monetization are an important driver for channels with problematic content to thrive [8, 43, 44].
Domain and Handle Reuse. Our work shares many similarities with work on residual trust across domains and namesquatting [45, 46]. Most related to our work, Mariconti et al. study the reuse of profile names on on Twitter [47]. Their study focuses on accounts that register handles that had been de-registered to capture the residual reputation and backlinks pointing to those handles [47]. However, while some of our findings, such as prevalence of issues and the fact that reuse is often used for questionable purposes [47], overlap, our work is substantially different, given that a repurposed channel (YouTube) carries all previous followers. The market for ready-made accounts has remained largely understudied, only receiving a brief overview by Chu et al. [8], and more recently a more in-depth characterization by Beluri et al. [7]. To our knowledge, our work is the first to explore how sold accounts are repurposed and to identify channel repurposing in the wild.
(a) A channel when listed for sale.
(b) Channel after reuse.
Figure 1: On the left, a channel listed for sale on Fameswap about entertaining facts with 1.19M subscribers. Videos are created predominantly with AI tools. This channel would later change to a news channel discussing political issues with no trace of its previous identity (on the right). For privacy, the channels’ handles and titles are blurred.
(a) Collection procedure for channels listed on Fameswap.
(b) Collection procedure for reference sample.
Figure 2: Data collection procedures. We collected Fameswap listings (1a) daily and a large sample of Social Blade channels (1b). We then scrape snapshots through time. On Fameswap, we scraped a channel every 3 days on average (2a). For Social Blade channels, due to the sample size, we collected a snapshot in January and later in March 2025 (2b). Finally, we obtain historical snapshots using the Wayback Machine (WM; see 3a and 3b).
3 Background
We provide background on the market we study and relate it to some of the earlier findings in the literature.
3.1 The Fameswap Market
Figure 3: Distribution of listing prices, subscriber counts, and price per 1,000 subscribers for Fameswap listings. Outliers (Q3 + 1.5X IQR) are hidden. Categories are self-reported at the time of listing. Categories are ordered by median price per 1,000 subscribers.
Fameswap is an online marketplace for buying and selling accounts on YouTube, TikTok, Instagram, Twitter/X, and websites. Recent measurements identify it as the largest such market by number of sellers [7]. The site resembles a standard e-commerce platform: each account has a dedicated listing with a description, price, seller reviews, and a content category (e.g., humor, sports). Buyers can bid, and the platform offers escrow, dispute resolution, and account verification—features similar to marketplaces like eBay.
Fameswap earns revenue from escrow fees (3% or a USD 50 minimum) and premium accounts, which offer advanced search, additional metrics, lower fees, and more. Payments are accepted via PayPal, wire transfers, and cryptocurrencies (e.g., BTC, ETH, USDC).
As of writing, 25,404 listings across all platforms are advertised for a combined total of USD 366.7M, claiming over 2.6B followers. For YouTube, we filtered for unique channel IDs, identifying 4,641 listings with 823M subscribers (confirmed via the YouTube API by March 31, 2025) and a total listed price of USD 160.4M. Chu et al. found 3,112 listings between 2019–2021, suggesting a nearly 9x growth [8]. The average listing price is USD 5,400 with 105,751 subscribers—comparable to Chu et al.’s findings. Like Beluri et al., we observed inflated pricing: 41 listings exceeded USD 100K and 16 exceeded USD 1M.
To contextualize these prices, we computed the cost per 1,000 subscribers—a common metric on engagement forums [18]. As shown in Figure 3, this varies widely by category. The median cost on Fameswap is substantially higher than the USD 16.52 per 1,000 subscribers found by Nevado-Catalán et al. [18]
Verifying sales remains difficult, as with other scraping-based estimates [48, 49]. During our study, we scraped 16,110 seller profiles—7,000 more than Beluri et al. [7]—identifying 2,930 escrow transactions and 684 reviews. Only 1,590 transactions disclosed sale amounts, totaling USD 1.16M. Though modest, this likely represents only a fraction of actual sales, which also occur via forums, Discord, Telegram, and private deals on Fameswap.  [^1] We suspect public sales skew toward smaller, cheaper channels, as the average sale price (USD 731) is far below the average listing price (USD 5,400).
3.1.1 Account Ownership Verification
Fameswap verifies account ownership by providing sellers with a unique randomized string that should be placed in the advertised social media profile. For example, a seller receives the string aSgc5H3s and places it temporarily in their YouTube channel description, allowing Fameswap to verify the ownership. This process is similar to using TXT records to verify domains.
3.1.2 Other Social Media Account Markets
Chu et al.[8] identified five marketplaces for social media accounts: SWAPD[50], Accs-Market.com [51], Trustiu (now inactive), ViralAccounts [52], and Fameswap [53]. We compare primarily with Chu et al.’s work, as Beluri et al.’s study became public after our data collection began and is referenced retrospectively. We also identified additional venues, including OGUser (similar to SWAPD) and Telegram channels.
Of these, only Fameswap and Accs-Market offer e-commerce-style interfaces, with listings, escrow, and reputation systems. ViralAccounts operates as a broker, and SWAPD functions as a private forum, limiting large-scale analysis. At study onset (October 2024), Accs-Market was inaccessible (HTTP 500), and archived pages returned HTTP 301 errors, preventing further analysis. We therefore focused on Fameswap, given its accessibility, scale, and apparent prominence.
3.2 YouTube Handles and IDs
In October, 2022, YouTube introduced handles as a way to uniquely identify accounts [54]. Different from channel titles, handles (“@SomeChannel”) are unique and can be accessed through “www.youtube.com/@SomeChannel.” Handles were rolled out gradually, but by the end of 2023 most accounts would have chosen or received a unique handle for their account. Handles resolve to a unique channel ID, allowing a channel to change its handle but keep pointing to the same channel. YouTube’s channel ID is a string that begins with “UC”, followed by 22 characters (letters, numbers, dashes, and underscores). Channel IDs can also be used to access a channel in the form: “www.youtube.com/channel/UC…”. Importantly, while an account can change its handle, its channel ID cannot change, which allows us to monitor accounts over time.
4 Methods
We measure the repurposing of YouTube accounts and their participation in disseminating problematic content through the lens of sold accounts extracted from a social media marketplace (Fameswap) and by observing handle changes and channel repurposing in the wild (from Social Blade). Specifically, we structure our study in three parts. First, we define, characterize, and estimate the prevalence of channel repurposing (Section 5). Second, we qualitatively annotate channels to identify potentially problematic content based on YouTube’s guidelines and prior work (Section 6). Third, we describe our framework for deriving statistically valid estimates using textual annotations and quantitatively test what features are indicative of repurposing (Section 7). An end-to-end view is later provided in Figure 9.
This section discusses our data collection practices and each sample used in the study. The following sections describe each experiment in turn.
4.1 Data Collection and Sources
4.1.1 YouTube Channels For Sale on Fameswap
As illustrated in Figure 2(a), we conducted daily scrapes of Fameswap from October 26, 2024, to March 31, 2025. The Fameswap interface provides a paginated list of all account listings. We scraped all historical listings. This set includes all listings not deleted or hidden from the website. On October 2024, Fameswap began displaying channel IDs with YouTube listings (whereas before they only disclosed the channel title). Together with our Fameswap scrapes, we began conducting regular scrapes of all YouTube channels advertised across Fameswap listings. We scraped channels every 3 days on average. As of March 31, 2025, we collected 4,641 YouTube channel IDs advertised for sale, each with about eight observations per month. For each channel we collected from Fameswap, we scraped their YouTube channel using the YouTube Data API v3.
4.1.2 Collecting a Large Sample of YouTube through Social Blade
To estimate the prevalence of channel repurposing and identify repurposing indicators beyond marketplaces, we aimed to observe this phenomenon “in the wild.” Randomly sampling YouTube is difficult due to the lack of a centralized channel directory and method to enumerate channel IDs [55]. Horta Ribeiro and West address this by sampling from Social Blade [56], an analytics site that has indexed over 68M channels in the past 17 years [57]. Social Blade is a widely used reference for creators [58] and has informed multiple YouTube studies [59, 60]. Building on this method, we sampled a large set of channels from Social Blade using a sample from Horta Ribeiro and West. While Social Blade’s crawling process is not public, it likely favors more popular channels. This, however, is acceptable given that our focus is content creators with larger audiences.
We sampled 1.4M channels and conducted an initial scrape from December 23, 2024, to January 21, 2025 (Figure 2(b)). We collected metadata from 1,397,586 channels and 139M corresponding videos. A second scrape, from March 21–31, 2025, captured updated metadata and 20M new videos. Only 1,351,912 channels returned data, indicating a 3.3% deletion rate. We used yt-dlp, a common tool for archiving YouTube content [61, 62, 63], discussed further in Section 9.
4.1.3 Samples Considered in this Study
  • Set of channels for sale on Fameswap: All channels collected from Fameswap, $n$=4,461 from October 21 2024, to March 31 2025 (see Section 4.1.1).
  • Set of repurposed channels from Fameswap: Subset of repurposed Fameswap channels, $n$=1,084 (see Section 5.2.1).
  • Large random sample from Social Blade: All Social Blade channels, appearing both in January and March 2025, $n$=1,351,912 (see Section 4.1.2).
  • Set of repurposed channels from Social Blade ($>$1,000 subscribers: Subset of Social Blade channels that we classified as repurposed, $n$=1,040 (see Section 5.2.1.
  • Random subsample from Social Blade (Baseline): Random subset of channels drawn from Social Blade with more than $>$1,000 subscribers, $n$=3000.
4.1.4 Historical View of YouTube Channels
To obtain a historical view of changes, we leveraged snapshots from the Wayback Machine [64]. The Wayback Machine is a digital archive initiative by the Internet Archive; it allows users to go “back in time” to see how websites looked in the past. The Wayback Machine allows users to capture pages for archival purposes [65]. The frequency of snapshots varies per website. More popular websites (e.g., higher-ranked by Alexa, higher number of inbound links, etc.) will be crawled more often [65]. Given a sample of YouTube channels, we attempt to obtain monthly snapshots from October 2022 to March 2025.
5 Repurposed Channels
Our first goal is to detect channels that have been repurposed, such as the example shown in Figure 1. In this figure, we observe how a channel that was posting primarily entertainment facts then becomes a channel that shares political news.
We first apply a qualitative approach to define what constitutes a channel being repurposed. We use these qualitative insights to develop an LLM prompt which we can then use to scale our annotations. By relying on an expertly coded subset, we can estimate the prevalence and error rate. Lastly, we explore the time it takes for a channel to be repurposed and a channel’s subscriber count after a channel has been repurposed.
5.1 Defining Repurposed Channels
To define what channel repurposing is, we followed a three-stage process. First, the lead author manually monitored a random sample of 104 channels (10%) that changed handles after being listed for sale. Each channel was visited weekly over a month to observe changes in content, metadata, and activity. We created standardized weekly snapshots capturing each channel’s videos, thumbnails, titles, timestamps, and descriptions (Appendix 10.1). Second, using open coding, two researchers independently analyzed a new random sample of 54 channels (5%) using these documents. Each was prompted: “Was the channel repurposed?” Coders discussed their annotations to identify edge cases and refine definitions. Disagreements arose around dormant channels, subtle content shifts, and stylistic similarities with different themes (e.g., AI-generated videos with varying topics). This coding was exploratory; we did not compute agreement scores. Coders prioritized changes in handles and titles, using descriptions and video content as secondary evidence. Minor changes (e.g., @HealthReporter to @TheHealthReporter) were not considered repurposed. Significant discrepancies (e.g., @HealthReporter to @247News) prompted checks for overlap in topics, URLs, or referenced entities. If no continuity was found, the channel was flagged as repurposed. A substantial shift in video content (e.g., religious to crypto content in another language) served as a confirmatory signal but was not sufficient on its own. In the third step, the lead author and an external coder with a large social media following reviewed another 54 channels (5%) as a sanity check to refine the working definition.
Definition: Repurposed Channel We consider a channel to have been repurposed from time $t$ to $t+1$ (two observations in time) if there is no perceivable association between the channel’s new handle and new title from its prior handle and title. Further, the channel description must not contain any references to its previous identity (e.g., “this channel was previously named X”), nor any overlapping text, nor any pointers to the same resources (e.g., a URL pointing to a social media account mentioned at time $t$).
Figure 4: Time (days), since a channel is listed for sale, until it changes its handle, title, and description (cumulative).
We created an LLM prompt based on our final definition (Appendix 10.1). Following, best practices suggested by Ziems et al. [66]. Using a prompt based on this codebook, we annotated Fameswap channels that had a changed handle using gpt-4o-2024-11-20 with temperature zero and top $p$ of one. Although there is still a lack of consensus on the optimal parameter choice for LLMs, recent work recognizes these parameters as the current standard [67]. We describe the validation in Section 7, alongside the framework we employ to obtain statistically valid estimates using text annotations.
5.1.1 Time Until Repurpose
By annotating Fameswap channels, we identified 1,084 channels (23%) (boasting 220M+ subscribers) as repurposed. For each channel, we computed how long it took them to change their handle, title, and description. As seen in Figure 4, the median time for a handle change was 28.5 days, followed by 29.3 days for a change in the handle and title, and 30 days for a change in handle, title, and description. As described in Section 4, we have an average of one observation every three days. Interestingly, handle, title, and description changes do not necessarily occur at the same time, but instead over a period of 36 hours (1.5 days), a fact that could inform an anomaly detection system. These results can be interpreted as a proxy for the median time of a sale.
5.1.2 Impact of Repurposing on Subscribers
To assess the impact of repurposing on subscriber count, we aligned all time series so that $t=0$ marks the week of the channel’s change (Figure 5). We tracked subscriber change (%) over the four weeks before and 12 weeks after this point. Most channels show slight but steady growth prior to the change, likely due to residual recommendation rather than active content updates. After repurposing, subscriber counts increase on average—often substantially. While some unsubscribes may be masked by new subscribers, the consistent growth suggests most users remain unaware of the change and stay subscribed.
5.2 Repurposed Channels in the Wild
Standard NLP approaches (e.g., edit distances, $n$-grams) struggle to detect meaningful change in handles and titles because they lack semantic understanding [68]. A significant transformation may be semantically stable (e.g., @TheLevenshtein to @EditDistance). LLMs are well-suited for this classification task (known as entity matching) given their semantic understanding of text, but also because of their memory [69]. Because they have been trained on large text corpora, LLMs can identify connections between entities a human annotator may miss, such as in the example above [69].
5.2.1 LLM and Expert Annotations of Channel Repurposing
From our sample of 1.4M YouTube channels, we detected 10,200 (0.73%) channels that changed their handle between January and March 2025. These handle changes do not necessarily mean a channel was repurposed. To identify channels that meet our repurpose definition (see Section 5.1), we use an LLM (prompt in Appendix 10.1) to classify each paired observation of channels. To validate the LLM’s annotation, we randomly sampled 10% of the channels ($n$=1,020) and manually annotated each pair of observations according to our definition. The false positive and negative rates are 4.9% and 3.8%, respectively.
Figure 5: Subscriber growth after handle change ($t$=0) for sold channels. Percentage growth is relative to the number of subscribers at $t$=0.
5.3 Prevalence Estimation
The LLM repurpose classifier achieved an accuracy between 94.2% and 96.6% (95% CI), resulting in an estimated 3,384–3,456 repurposed accounts. These results indicate that, from January 2025 to March 2025, approximately 0.24-0.25% of channels in the Social Blade population (68M+ channels) were repurposed. Of these repurposed channels, 1,074 had more than 1,000 followers. In total, these repurposed accounts have a total audience of 43,975,420 subscribers, a meaningful audience on YouTube during this period. We summarize the descriptive statistics in Table I.
5.4 Repurposed Channels Before Observation
As described in Section 4.1.4, we queried the archive for snapshots taken after YouTube introduced handles (October 2022) for all Fameswap-listed channels (n=4,461), including those not yet repurposed, and all repurposed Social Blade channels (n=1,040).
For Fameswap, we found snapshots for 819 channels (17.6%). Of these, 332 (40.5%) had a different handle than the one listed at the time of sale, and 71 (9.5%) showed more than one handle change across archived versions. Among the repurposed Social Blade channels, 289 (27.8%) had snapshots; 116 (40.1%) had previously used different handles, and 29 (10.0%) had changed more than once. Due to the absence of additional metadata (e.g., channel descriptions), we could not assess whether these changes were substantial or cosmetic.
The frequency of handle changes may be meaningful. Prior work by Mariconti et al. [47] found that Twitter accounts with multiple profile name changes were more likely to engage in misbehavior—a pattern that may hold for repurposed YouTube channels as well.
6 Content Analysis
Next, we explore the content disseminated through repurposed channels. Reusing a channel allows a user to distribute content to an existing audience or potentially increase the chances that the recommendation algorithm amplifies new content [70, 71]. Our goal is to better characterize the actors’ motivations through the lens of the content they produce after repurposing a channel. To this end, we begin by identifying relevant types of potentially problematic content, based on YouTube’s community guidelines, and define a set of codes. Using these codes, we annotate the videos on repurposed YouTube channels. Lastly, we explore what types of content repurposed channels were creating prior to changing and estimate their survivability.
6.1 Codebook Development and Annotation
We use an LLM to extract potentially problematic content in a channel. To narrow down the types of content for which we should annotate, we first conducted a qualitative analyzed channels and their contents.
6.1.1 Content Categories
To develop our category annotation codebook, we began with a deductive approach, selecting a set of restricted themes (or categories) from YouTube’s community guidelines [72]. However, YouTube’s content restrictions are often broad, cover a wide range of topics, and definitions often lack specificity [73]. To overcome this, we opted to empirically choose the topics we annotate and based on their presence in a smaller random subset. This approach lets us develop an intuition for the types of content disseminated and how they are presented, which will later be necessary to produce few-shot examples for LLM annotation. To begin this analysis, the lead author reviewed the videos of 104 repurposed Fameswap channels (10%), flagging videos that had content covered by community guidelines.
Annotating for kids content was straightforward based on YouTube’s definition [74], as well as content that may infringe copyright [75]. Similarly, based on the misinformation guidelines [76, 77, 78], we created three codes covering political, medical, and news content. Although news and political content overlap (i.e., politcal news), non-news political content, and non-political news content also exist and are important to distinguish [79]. We do not attempt to verify whether videos’ content constitute misinformation or disinformation—the latter define by its intentionality. We address this limitation in Section 8.1. In addition to political and news content, we noticed a substantial number of religious videos. Although religion is not a topic captured in community guidelines, we opted to include it as a code given its frequent co-occurrence with geopolitical content and frequently observed content involving religious leaders with political appointments. Past work has also found religious videos to frequently co-occur with extremist and hateful content [80, 81].
Figure 6: Topic correlation matrix. Values represent the Pearson correlation between variables.
On the financial side, we created a code for Gambling content, which is directly covered by the guidelines [82]. However, we note a substantial amount of content related to making money online through a variety of ways: pay-per-click websites, dropshipping, trading stocks, cryptocurrencies, get-rich-quick schemes, and ironically, courses to grow and monetize social media accounts. These topics are only partially covered by YouTube’s guidelines (e.g., get-rich-quick schemes, fraud, broadly “spam”, and broadly “scams”) [83]. Cryptocurrency-related content was particularly prevalent. Informed by research on cryptocurrency scams and their potential to harm users financially [84, 34], we create a cryptocurrency code. We capture the rest of the aforementioned content under money-making content.
Figure 7: Topic presence for reused Fameswap channels. Topics on the left are self-reported at the time of listing. Topics on the right are labels assigned during our topic detection. Flows are proportional and one-to-many. That is, if a channel has several topics, flows are drawn to all destination topics proportionally.
Using this codebook, two authors with extensive experience with the YouTube ecosystem independently coded 54 repurposed Fameswap channels (5% of reused channels). Each researcher was given a summary text document containing all of the channel’s observation snapshots. We provide an example of this document in Appendix 10.1. After the first round of coding, the coders discussed the procedure and tweaked the codebook to address the shortcomings. We also extracted additional codes from the open-ended field if they matched community guidelines violations or were potentially harmful and not yet captured by existing codes. In this additional round, we detected and included a category for extremist and manosphere content [56] through this procedure.
We repeated an additional coding round between a member of the research team and a member external to the research team with extensive knowledge about the creation of social media content. They coded a new randomly sampled but non-overlapping 5% of the dataset. The goal was to identify ambiguity in the existing definitions. Together with this external member, we derived our final codebook. We then converted our final codebook to a prompt for annotation. These artifacts can be found in Appendix II and  10.1, respectively. Note, we did not compute agreement during these steps since the goal of these steps was to create a codebook, not to use the codes for downstream analyses. We address the statisical soundness of our annotations in Section 7.
Figure 8: Survivability curve for Fameswap channels with both potentially problematic content detected and without. “Death” event is channel removal. Bands represent 95% CI.
6.1.2 LLM Annotations
Using a prompt based on this codebook, we annotated three sets of channels using gpt-4o-2024-11-20 with temperature zero and top $p$ of one. We annotated: 1) the set of Fameswap channels that changed handles ($n$=1,084); 2) the set of repurposed channels we identified in our reference sample with over 1,000 followers ($n$=1,074), and; 3) a random set of channels without handle changes drawn from the reference sample as a baseline ($n=$3,000). All sets are described in Section 4.1.3.
We feed the LLM up to 50K tokens per channel, combining the channel description with video titles and descriptions. Most channels fit this limit. If a channel exceeds it, we (i) de-duplicate strings with $>$ 0.9 edit similarity to capture more diverse content, then (ii) randomly sample the remaining text until the total is $\leq$50K tokens. We do this to capture a greater diversity of information, given that some channels post many videos with the same or similar titles. We cap input at 50K (well below GPT-4o’s 128 K window) because longer contexts seem to degrade classification accuracy [85].
6.2 Presence of Problematic Content
To better understand how problematic content manifests in repurposed channels, we examine three aspects. First, we analyze how different content types co-occur by computing correlations between annotated topics. This helps reveal common content pairings, such as between financial and ideological themes. Second, we trace how channels transition from their original categories to new potentially problematic categories after being repurposed. Using self-reported Fameswap categories and annotations after a channel has been repurposed, we highlight shifts from innocuous to concerning content types. Third, we assess whether such transitions are met with platform-level enforcement by measuring channel suspension rates over time. These analyses shed light on the nature, evolution, and consequences of problematic content in repurposed channels.
Figure 9: End-to-end experimental procedure, including samples, annotations, featurization, and regression.
TABLE I: Regression results and descriptives for Fameswap and Socialblade models.
(a) Logistic Regression Results
Fameswap Socialblade
Coefficient (SE) Coefficient (SE)
(Intercept) $\phantom{-}0.8332^{***}$ $(0.0982)$ $\phantom{-}0.3401^{***}$ $(0.0824)$
[…]
Comm. Ct. (Mean) $-0.0006^{*}\phantom{**}$ $(0.0003)$ $-0.0005\phantom{{}^{***}}$ $(0.0004)$
[…]
T. Btw. Vids. (SD) $\phantom{-}0.0049^{***}$ $(0.0013)$ $\phantom{-}0.0022^{***}$ $(0.0003)$
[…]
T. 1st to Oldest Vid. $\phantom{-}0.0243^{***}$ $(0.0043)$ $\phantom{-}0.0224^{***}$ $(0.0027)$
Oldest Vid. Age $-0.0702^{***}$ $(0.0038)$ $-0.0375^{***}$ $(0.0025)$
Subs. Count $\phantom{-}0.0000^{*}\phantom{**}$ $(0.0000)$ $\phantom{-}0.0000^{*}\phantom{**}$ $(0.0000)$
Total Videos $-0.0001^{*}\phantom{**}$ $(0.0001)$ $-0.0002\phantom{{}^{***}}$ $(0.0001)$
Content Features
[…]
AI-generated $\phantom{-}1.0935^{***}$ $(0.2057)$ $\phantom{-}0.6283^{***}$ $(0.2000)$
[…]
Religious $\phantom{-}0.9767^{***}$ $(0.1608)$ $\phantom{-}0.3910^{***}$ $(0.1187)$
News $\phantom{-}0.6627^{***}$ $(0.1918)$ $\phantom{-}0.4305^{***}$ $(0.1356)$
Medical/Health $-0.4690^{**}\phantom{*}$ $(0.1866)$ $-0.3025^{**}\phantom{*}$ $(0.1300)$
Cryptocurrency $\phantom{-}1.6575^{***}$ $(0.2491)$ $\phantom{-}0.5112^{*}\phantom{**}$ $(0.2424)$
[…]
Kids $\phantom{-}0.7105^{***}$ $(0.2130)$ $\phantom{-}0.3357^{**}\phantom{*}$ $(0.1341)$
P. Copyright Infr. $\phantom{-}0.4544^{***}$ $(0.1444)$ $\phantom{-}0.2826^{**}\phantom{*}$ $(0.0943)$
[…]
${}^{***}p<0.001$, ${}^{**}p<0.01$, ${}^{*}p<0.05$
(b) Descriptive Statistics
Fameswap (n=1,084) Socialblade (n=1,074) Baseline (n=3,000)
Mean SD Mean SD Mean SD
155,519 1,104,304 119,214 1,016,701 211,258 1,085,125
448,550 1,552,437 195,928 1,100,404 442,672 2,061,892
5,838 36,336 1,120 8,357 1,657 9,986
14,959 63,326 1,761 9,201 3,290 15,435
84.65 1,021.2 45.09 270.5 88.28 400.0
246.24 2,144.2 73.72 303.6 166.17 752.0
862.91 4,215.2 518.60 1,207.4 616.16 1,113.2
1,803.2 19,831 580.82 1,489.1 803.19 8,011.8
17.73 58.02 119.73 328.23 72.11 213.58
41.14 106.6 206.89 381.47 115.11 200.46
5.53 8.82 3.01 5.59 3.65 18.43
4.86 10.09 2.37 6.22 3.08 20.60
17.65 22.82 57.72 64.77 84.15 62.78
20.06 23.08 65.78 68.18 132.43 56.53
202,981 1,013,686 40,952 169,278 103,427 1,002,724
158.01 474.26 127.09 806.19 373.66 1,365.7
Content Features
0.347 0.476 0.372 0.483 0.432 0.495
0.105 0.307 0.066 0.248 0.022 0.147
0.210 0.408 0.171 0.377 0.187 0.390
0.203 0.403 0.194 0.396 0.154 0.361
0.176 0.381 0.154 0.361 0.129 0.335
0.090 0.286 0.121 0.327 0.153 0.360
0.111 0.315 0.039 0.194 0.014 0.118
0.053 0.225 0.021 0.143 0.017 0.129
0.160 0.367 0.113 0.316 0.073 0.260
0.047 0.212 0.062 0.241 0.075 0.264
0.251 0.434 0.345 0.476 0.426 0.495
0.036 0.186 0.011 0.106 0.006 0.078
6.2.1 Topic Correlation
We expected specific topics to be correlated, such as channels containing content about investing in stocks (“money” topic) and cryptocurrency. To investigate this, we compute the correlation matrix for all labels, shown in Figure 6. A topic correlation matrix for binary (0/1) topics shows the pairwise association between topics, specifically how often one topic is related to another across observations. A $\phi$ coefficient, which is shown for each pairwise relationship, of one would indicate that topics always co-occur, zero that they are independent, and negative one that they are anti-correlated. Expectedly, the highest correlations are between political and news content and money and cryptocurrencies. We observe a weak correlation between medical, political, and news content. We also observe a weak correlation between manosphere, political, and news content. In particular, we find a weak correlation between gambling content and non-YouTube monetization. This indicates that gambling-related channels have more information on outbound links to purchase products or services, similar to channels with cryptocurrency content. In general, we find that content that may infringe copyright and AI-generated content appears throughout the board, independent of other topics, having only a slight association with political, money-related, and child content. Lastly, while there is no strong correlation, we find that kids content co-occurs with potentially problematic categories. As noted by past work, this can particularly harmful for young audiences [86].
6.2.2 From Innocuous to Problematic
When listing a channel for sale on Fameswap, users choose a category that best fits the channel’s content. Choosing a category improves the searchability of the channel in the marketplace. As observed in Figure 3, there are 18 categories that cover a wide range of topics. More than 88% of the listings have categories. We manually verified the validity of these self-assigned categories and found them to be largely accurate. Using these self-assigned categories, we investigated channel transitions into potentially problematic categories, as shown in Figure 7. We find that 404 (37%) of repurposed channels later displayed problematic content in their channels. Of these, 25% had ideological content and 12% had financial content. In particular, the origin category (from Fameswap) had little to do with the types of problematic content that would appear in the channel. We see that every innocuous category can ultimately disseminate problematic content.
However, not all channels necessarily transitioned into problematic categories. During our observation period, there were many channels in which we detected no objectionable content even after being repurposed. For example, a sports page became a musician’s page, and an entertainment page later became an influencer’s travel vlog. At the same time, some repurposed channels would later become channels associated with companies or brands. Although they were not posting objectionable content, it raises the question of whether there is a need for more transparency, given that people may rely on metrics such as subscriber counts to judge the authenticity of a company.
6.2.3 Rate of Suspension
iven the granularity of our observations for Fameswap channels, we estimate the rate at which YouTube suspends these accounts. To do so, we compute the survival function for two groups, repurposed Fameswap channels and those not yet repurposed, using the Kaplan-Meier estimator. As shown in Figure 8, all Fameswap channels exhibit an average survivability of approximately 85% after 200 days (95% CI: 79%–87.5%). Based on a log-rank test, there is no significant difference between repurposed channels and those that have yet to be sold or repurposed. Using our Social Blade scrapes, we computed the percentage of deleted or suspended channels between January and March 2025. After 90 days, we observed that 3.3% of the channels were deleted. That is, 96.7% channels from the larger Social Blade sample ($\sim$1.4M) survived. These results indicate that Fameswap channels have slightly lower survivability ($\sim$91at 90 days). Given that 37% of the repurposed channels displayed some potentially problematic content, we expected a lower survival. However, these results indicate that repurposed channels and channels for sale generally avoid penalties.
7 Indicators of Channel Repurposing and Prevalence
We next identify the indicators that correlate with repurposed channels in our two channel samples: channels originating from Fameswap and those collected in the wild from Social Blade. To do this, we employ a logistic regression model with the same set of features and compare each sample against a baseline. We confirm that many topics discussed in Section 6 are commonly associated with channel repurposing. In addition, we identify quantitative features, such as video likes and posting behavior, that also predict channel repurposing. The goal of these experiments is to quantitatively understand what features, if any, are commonly associated with channel repurposing. Additionally, we want to test whether the potentially problematic topics that we annotated appear at a significantly higher rate across repurposed channels compared to regular channels.
7.1 Regression Model
We use a logistic regression model to examine the relationship between channel characteristics and channel repurposing, modeling the log-odds of the outcome as a linear combination of the predictors, in our case:
$\log\left(\frac{p}{1-p}\right)=\beta_{0}+\beta_{1}F_{1}+\cdots+\beta_{j}F_{j}\\ +\beta_{j+1}T_{1}+\cdots+\beta_{j+k}T_{k}\ ,$ (1)
where $\beta_{0}$ is the intercept (log-odds outcome when all predictors are zero), $F_{i}$ represents quantitative features extracted from each channel, and $T_{i}$ represents an indicator for each topic. In this model, a one-unit change in a covariate $X_{i}$ changes the log-odds outcome by $\beta_{i}$; equivalently, the odds change by a factor of $e^{\beta_{i}}$.
7.1.1 Quantitative Feature Selection
We take the mean and standard deviation of all observed video views, likes, and comments. We also compute the mean and standard deviation of the number of videos posted weekly, video lengths (seconds), and the time between each video posted (days). Finally, we compute the time between the first and last video the channel has posted (months) and the oldest video the channel has, which would typically mark when the channel first started uploading content. Finally, we take the total number of videos. Table I presents the descriptive statistics for each sample. Video deletions are likely an important feature. However, because we do not have frequent snapshots of Social Blade channels (given the sample size), we miss all uploads and deletions that take place between our two Social Blade snapshots. However, we argue that our current results are robust to the omission of this covariate given that the time between videos may still capture this signal.
7.1.2 Using Text Annotations with Valid Estimates
We combined small-scale qualitative annotations with large-scale LLM predictions, using the latter as surrogate classifiers to enable scalable analysis. LLMs have significantly advanced computational social science [66], with growing evidence supporting their reliability across disciplines [87, 66] and platforms like YouTube [88, 79]. However, using LLM-generated labels in downstream analyses introduces measurement error—mismatch between predicted and true labels [89]. To address this, we follow Egami et al.’s framework for statistically valid inference with LLMs [89], using the Design-based Supervised Learning (DSL) framework. The process involves: (1) predicting labels with an LLM, (2) sampling a subset for expert annotation, and (3) combining predictions and gold-standard labels in DSL regression. This method requires that expert-coded samples have known, non-zero sampling probabilities. In our case, we created two such datasets: one in which we annotated whether a channel was repurposed or not (Section 5.2.1) and one where we extracted topics (Section 6.1.2).
7.1.3 Expert Topic Annotations
We create an expert-annotated (“gold-standard” ) dataset of topic extraction, by annotating a sample of 260 channels (5% out of 5,158) randomly drawn from the three groups for which we extracted topics with an LLM: repurposed Fameswap channels, repurposed Social Blade channels, and a baseline described above. These channels had a combined total of 14,354 videos, which a coder manually annotated. To create a valid expert-annotated dataset for channel repurpose, we leverage the Social Blade sample we annotated in Section 5.2.1 and annotate a 10% sample of repurposed Fameswap channels.
7.2 Regression Results
We conduct two logistic regressions, first we model repurposed channels from Fameswap against a baseline drawn from Social Blade using the R package,  DSL [90]. We report the coefficients in Table I.
Takeaways from Quantitative Features: Channels whose oldest video is more recent are more likely to be repurposed, as are those with a longer interval between oldest and most recent uploads. So, buyers seemingly favor channels that look established (i.e., showing a sizeable upload time span), but those are not necessarily truly long-standing. Indeed, a genuinely mature channel would feature both an old first upload and a long production span.
Irregular upload intervals (high standard deviation among “time-between-videos”) and modestly larger subscriber bases both predict repurposing, implying that erratic schedules and an existing audience are typical markers of repurposed channels. Mean views/likes do not have statistically significant effect on predictions. Lower comment counts are, surprisingly, predictors of channel repurposing. A plausible hypothesis is that repurposed channels may disable or limit comments to avoid scrutiny. Subscriber count is a positive but modest predictor of channel reuse: repurposing seems less about absolute reach and more about fast content throughput in lucrative niches.
Takeaways from Qualitative Features: Repurpose odds are primarily content-driven. AI-generated content ($\beta$=0.6–1.1) is the strongest predictor of repurposing, probably because automated pipelines scale cheaply after a channel is acquired. However, we rely on disclosure to tag channels, as described in Section 6, so so the coefficient may also reflect self-disclosure bias. Cryptocurrency shows the next-largest effect ($\beta$=1.7 on Fameswap), followed by gambling. We hypothesize that these categories are prone to reuse because of their monetization potential. In line with previous work, we find a significant amount of off-platform monetization [43, 8], albeit only for Fameswap. Surprisingly to us, medical/health is negatively associated, suggesting either stricter platform scrutiny or lack of prominent health-related narratives. Religious content and potentially copyright-infringing content are associated with repurposing. A possible explanation could be that these types of content are used to grow channels due to their appeal (e.g., free TV shows), similar to kids content, and we are identifying trace videos prior to deletion.
8 Discussion and Conclusions
Our results suggest that channel repurposing involves sold channels (such as those advertised on Fameswap), as well as channels that may not have been for sale (such as most of those found through Social Blade). In either case, we find that channels with large numbers of subscribers—220M+ for Fameswap and 43M+ for Social Blade—completely repurposed their channels, erasing any perceivable association with their prior identity. Across most Fameswap accounts, channel repurposing seemed to go unnoticed, as indicated by the subsequent growth in subscribers. The median time between a channel being listed for sale and being repurposed was 30 days. However, some took as long as 200+ days (RQ1).
On Fameswap, 37% of channels posted potentially problematic ideological and financial content. Before these accounts were repurposed, they belonged to various innocuous content categories, ranging from humorous content to sports and celebrity gossip. Figure 1 illustrates an example of a typical problematic transition, in which a channel that previously shared ‘interesting facts’ is sold and repurposed into a news channel featuring contentious geopolitical actors: Russia, Ukraine, Zelenskyy, etc. Due to space constraints, we cannot discuss all remarkable cases individually, and instead, quantitatively show that potentially problematic content is significantly related to channel repurposing (RQ2).
By leveraging a large sample of YouTube channels, from a random Social Blade sample, we estimated that 0.24-0.25% of channels were repurposed between January and March 2025. These results imply that there were another 16,000+ repurposed channels (out of the 68M captured by Social Blade) that we did not capture, likely with tens or hundreds of millions of subscribers (RQ3).
Lastly, engagement metrics offer little indication that a channel has been or will be repurposed. Instead, video upload behaviors (i.e., time between videos) and potential gaps in their video history are better indicators for potential repurposing. However, these indicators may not be detectable or readily interpretable by everyday users (RQ4).
Our results suggest that social media audiences are becoming more commoditized (i.e., have less differentiation, have widespread availability, and compete mainly on price). From the lens of transaction cost economics, social media account markets are evolving from forums (unassisted markets) into assisted markets, with lower barriers to entry and increased liquidity [91]. This transition typically makes outsourcing more attractive and, as a result, fosters specialization, a pattern that frequently emerges across cybercriminal endeavors [92, 93, 94]. In our context, this may signal the emergence of actors who specialize in harvesting organic audiences.
We do not claim nor expect these markets to be free of fraud. Many vendors may use artificial engagement to inflate their numbers and sell accounts at a premium. However, these markets are increasingly becoming more transparent, allowing prospective buyers to access more information. This continuous evolution, coupled with the apparent success of these markets, seems to indicate that a substantial portion of buyers find these markets useful.
We found evidence that channels may be repurposed into disinformation outlets. Additionally, generative artificial intelligence (GenAI) seems to play an important role in the automation of content creation, as we found evidence of GenAI-created satirical political content. In general, repurposed channels show a high correlation with AI-generated content. Putting misinformation consumption in terms of supply and demand, scholars have argued that AI does not increase the demand for misinformation [95]. However, our results add nuance to these claims. By facilitating the cultivation of audiences that can then be served mis/disinformation, GenAI may be helping increase the demand for disinformation indirectly. Furthermore, we hypothesize that GenAI enables actors to more easily produce diverse content across niches and languages, facilitating the cultivation of specific audience demographics—an asset that can be strategically leveraged when a channel is repurposed to target vulnerable populations.
There are many legitimate reasons to change handles, titles, and descriptions. Likewise, benign users may want to change their handles without drawing unwanted attention. However, current changes are imperceptible to subscribers, which leads to potential abuse. We do not suggest that platforms such as YouTube start monitoring and banning accounts from second-hand social media account marketplaces or seek to deplatform these marketplaces. These interventions will likely only succeed in making nefarious activity harder to measure [96]. Instead, a simpler start could be to explicitly call out changes in channels above a certain number of subscribers, with varying degrees of notification depending on the channel size and the significance of the change. Ultimately, improving visibility into major channel changes rather than suppressing second-hand marketplaces themselves offer a path forward, preserving user awareness while maintaining the ability to monitor this evolving ecosystem.
8.1 Limitations and Future Work
A main limitation of our study is that we rely on video metadata to infer topics, without analyzing the video content itself. Thus, when labeling channels as political, medical, or news-related, we do not assess whether the content is actually harmful or mis/disinformation. Fact-checking video content remains underexplored, with most work focused on news articles [97]. Assessing harm is also inherently subjective. While some cryptocurrency or money-themed channels promoted scams, we cannot generalize this finding to the entire financial category. Still, identifying channels with potential for harm is a first step. Future work could analyze transcripts, video frames, or external URLs to refine these labels and estimate the volume of problematic content rather than relying on binary classifications.
Although Fameswap is likely a key player in the second-hand social media account ecosystem, further work is needed to assess its influence. For instance, the extent of overlap across marketplaces or between marketplaces and forums is unknown. This motivated our prevalence estimation, which avoids limiting analysis to marketplace samples. As discussed in Section 3, Fameswap verifies ownership using unique strings. Future work could develop tools to detect these in the wild. Finally, our study focuses only on YouTube. While we expect our findings to generalize to platforms like TikTok, Twitch, or Twitter/X, this remains to be tested.
9 Ethics
Purchasing a Fameswap membership raises ethical concerns, as it supports a service tied to potentially questionable practices. However, Fameswap is not illegal, nor does YouTube explicitly prohibit channel resale [98]. In the absence of alternatives, and given the modest cost (US\$ 348 total), our active approach aligns with research precedent. Prior work has involved similar or greater expenditures: Karami et al. [99] spent USD 140 on DDoS services; Levchenko et al. [100] and Kanich et al. [101] made over 250 purchases totaling several hundred dollars across unlicensed pharmaceutical outlets; Mirian et al. [102] spent USD 1,633 hiring hackers. All these numbers are in line with or exceed our own expenses in this project.
We used yt-dlp to collect metadata from YouTube. Although the RIAA issued a DMCA takedown in 2020 over its use for video downloads, the notice was later reversed [61]. Even then, metadata collection, our primary use case, was not the focus of the takedown. yt-dlp remains widely used in academic research [63]. To minimize platform impact, we prioritized existing datasets (obtained from Horta Ribeiro and West [55]) and employed conservative rate limits. We followed ethical data practices aligned with Beadle et al. [103]. (YouTube does not meet Beadle et al.’s definition, which is based on a list compiled by Wikipedia[104].) We limited our analysis to larger channels (those with $>1,000$ subscribers), which we consider to have a reduced expectation of privacy due to their public reach. All data was stored on a secure university server and accessed exclusively by the research team, unless otherwise noted.